OpenStudy (dreamie98):

Hello, can someone please explain me how to solve quadratics by factoring. I can't seem to get the right answer with this one; x^2 - 2x + 1 = 0 So this is how I solved it but its wrong; =(x-1)(x+1) =x-1 +1=0 +1 x=1 =x+1 -1=0 -1 x = -1 Apparently, the answer is supposed to be two positive ones and I'm getting a positive and a negative one T-T PLEASE HALP. PS: I really have no idea when its supposed to be a negative or a positive in the factoring process.

Nnesha (nnesha):

\[\huge\rm ax^2+bx+C=0\] when the leading coefficient is 1 find two numbers when you multiply them you should get product of `AC` (a =leading coefficient , c=constant term) and when you add/subtract them you should get middle term which is `b` (coefficient of x term) check your answer (x-1)(x+1) is -1 *1= C ?? -1+1 = b ??

Nnesha (nnesha):

correction -1 *1= A C ?? -1+1 = b ?? what two numbers would you multiply to get `1` but when you add same numbers you should get `-2`

OpenStudy (dreamie98):

-1 :o?

OpenStudy (dreamie98):

like -1 x -1 = 1 and -1 + -1 = -2 ?

Nnesha (nnesha):

yes right so it should be (x-1)(x-1 ) correcto ??

OpenStudy (dreamie98):

yuuus. BUT WAIT, thats the answer I got when I solved it with the formula, but when I did it by factoring i got +1 and -1 AND the answer sheet says its only +1 and +1. IM CONFUSED.

Nnesha (nnesha):

(x+1)(x-1) isn't correct Because -1 times 1 isn't equal to +1 -1+1 isn't equal to -2

Nnesha (nnesha):

the two numbers that will give you the product of AC and sum o b are -1 and -1

OpenStudy (dreamie98):

and when Im factoring the original quadratic which is x^2 -2x + 1 = 0 How do I know if its positive or negative? T-T BTW does your shortcut method work 100% with harder quadratics too ?

Nnesha (nnesha):

yes it always work if the leading coefficient is one its known as headphone method  okay how would you factor this x^2-2x+1= 0???